# Quantumness of gravitational cat states in correlated dephasing channels

### v1

@@ Abstract
Quantumness; gravitational cat states; correlated dephasing channels
@@ I Introduction
The study of gravcat states has attracted considerable attention in recent years [8, 9, 10, 11], which has been stimulated by advances in theoretical frameworks and experimental techniques. Some researchers have proposed various schemes to generate and observe gravcat states, utilizing tools from gravitational wave detection, quantum optics, and precision measurement techniques. These efforts not only deepen our knowledge of fundamental physics but also open avenues to explore the quantum nature of gravity in new ways. Some papers [12, 13, 14, 15] highlight theoretical proposals and experimental approaches towards understanding and realizing gravcat states.
@@ II Gravcat model and method
with $H_{\omega}=\omega(\sigma_{z}\otimes\mathbf{I}_{2}+\mathbf{I}_{2}\otimes\sigma_{z})/2$ and $H_{\gamma}=-\gamma(\sigma_{x}\otimes\sigma_{x})$, where $\sigma_{x}$ and $\sigma_{z}$ are Pauli operators and $\mathbf{I}_{2}$ is identity operator. Additionally, $\omega$ represents the energy gap between the ground state and the first excited state, and $\gamma=\frac{Gm^{2}}{2}\left(\frac{1}{x}-\frac{1}{x^{\prime}}\right)$ serves as a parameter controlling the strength of the gravitational interaction between states with masses $m$. Herein, $G$ denotes the universal gravitational constant, and $x$ ($x^{\prime}$) represents the distance between two masses when they are at the same (different) relative minimum.
@@
Due to the functional equation expressed in (2) and the structure of our Hamiltonian as given in Eq. (1), the Gibbs density matrix would be obtained as
@@ III Gravcats in correlated dephasing channel
where $E_{ij}=\sqrt{p_{i}p_{j}}\sigma_{i}\otimes\sigma_{j}$ are the Kraus operators with identity operator $\sigma_{0}=\mathbf{I}_{2}$ and three Pauli operators $\sigma_{1,2,3}=\sigma_{x,y,z}$. The probability distribution is defined by $\{p_{i}\}$ with $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$. Besides, $\{E_{ij}\}$ satisfy the completely positive and trace preserving condition as $\sum_{ij}E_{ij}=\mathbf{I}_{4}$.
@@ IV Quantum coherence of gravcats
V Fisher information of gravcats
Quantum Fisher information (QFI) originates from quantum metrology, focusing on measurement precision in the quantum domain [30, 31, 32, 33]. In quantum mechanics, Fisher information quantifies the information regarding a parameter contained in measurement outcomes, indicating how well a quantum state can be distinguished from nearby states concerning the measured parameter [34]. The local quantum Fisher information (LQFI), denoted by $\mathcal{F}$, builds upon QFI ($F$), representing the optimal LQFI when $H_{A}$ (a measurement operator) is applied to subsystem $A$ belonging to bipartite system $AB$. Therefore, LQFI can be expressed as follows [34]
@@ VII Comparative analysis of quantum resources
For $\mu=1$, as depicted in Fig. 8, all quantities remain constant over time, which indicates complete robustness. This finding underscores the critical role of classical correlations in preserving quantum resources. Conversely, for $\mu<1$, exponential decay is observed in all quantities. Lower values of $\mu$ correspond to more pronounced decay, highlighting the detrimental effect of reduced classical correlations on the preservation of quantum resources. In the extreme case of $\mu=0$, quantum coherence and LQFI asymptotically approach zero, while the Bell parameter converges to $2$, indicating the loss of non-classical correlations.

### v2

@@ Abstract
Quantumness; gravitational cat states; correlated dephasing channels; classical correlations
@@ I Introduction
Historically speaking, there has been ongoing discussion around a fundamental question related to this issue: how can we confirm if gravity must be treated as a quantized phenomenon or “why we need to quantise everything, including gravity” [8]? Furthermore, is there a universally applicable experimental method that can determine whether gravity operates on a quantum level [9, 10]? According to certain perspectives drawing from quantum information theory, it is posited that an interaction capable of generating entanglement between two systems must inherently possess quantum characteristics. Consequently, a significant indicator of quantum gravity would be the observation of entanglement between massive states brought about by gravitational interactions [11, 12]. Contrary to current tests that rely on detecting entanglement mediated by gravity, Lami et al. [13] have recently proposed a new approach focusing solely on coherent states. Interestingly, their method eliminates the need to generate extensively delocalized states of motion or detect entanglement, which does not occur at any stage of the process.
Hence, the study of gravcat states has attracted considerable attention in recent years [14, 15, 16, 17], which has been stimulated by advances in theoretical frameworks and experimental techniques. Some researchers have proposed various schemes to generate and observe gravcat states, utilizing tools from gravitational wave detection, quantum optics, and precision measurement techniques. These efforts not only deepen our knowledge of fundamental physics but also open avenues to explore the quantum nature of gravity in new ways. Some papers [18, 19, 20, 21] highlight theoretical proposals and experimental approaches towards understanding and realizing gravcat states.
@@ II Gravcat model and method
with $H_{\omega}=\omega(\sigma_{z}\otimes\mathbf{I}_{2}+\mathbf{I}_{2}\otimes\sigma_{z})/2$ and $H_{\gamma}=-\gamma(\sigma_{x}\otimes\sigma_{x})$, where $\sigma_{x}$ and $\sigma_{z}$ are Pauli operators and $\mathbf{I}_{2}$ is identity operator. Additionally, $\omega$ represents the energy gap between the ground state and the first excited state, and $\gamma=\frac{Gm^{2}}{2}\left(\frac{1}{x}-\frac{1}{x^{\prime}}\right)$ serves as a parameter controlling the strength of the gravitational interaction between states with masses $m$. Herein, $G$ denotes the universal gravitational constant, and $x$ ($x^{\prime}$) represents the distance between two masses when they are at the same (different) relative minimum (look at Fig. 1).
@@
Due to the functional equation expressed in (2) and the structure of our Hamiltonian as given in Eq. (1), the Gibbs density matrix would be obtained as [15, 16, 17]
@@ III Gravcats in correlated dephasing channel
where $E_{ij}=\sqrt{p_{i}p_{j}}\sigma_{i}\otimes\sigma_{j}$ are the Kraus operators with identity operator $\sigma_{0}=\mathbf{I}_{2}$ and three Pauli operators $\sigma_{1,2,3}=\sigma_{x,y,z}$. The probability distribution is defined by $\{p_{i}\}$ with $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$. Besides, $\{E_{ij}\}$ satisfy the completely positive and trace preserving condition as $\sum_{ij}E_{ij}^{\dagger}E_{ij}=\mathbf{I}_{4}$.
@@
Note that we have $\eta=1$ for a fully correlated dephasing channel, i.e. $\mu=1$. So, the density operator (9) will be time-independent in this particular case (gravcats).
@@
(a)Refer to caption(b)Refer to caption(c)
Figure 10: The $\mu$ dependence of quantum coherence (a), LQFI (b), and Bell non-locality (c) with fixed $f=0.8,0.6,0.4,0.2$, and 00 (from top to bottom).
@@ IV Quantum coherence of gravcats
V LQFI of gravcats
Quantum Fisher information (QFI) originates from quantum metrology, focusing on measurement precision in the quantum domain [36, 37, 38, 39]. In quantum mechanics, Fisher information quantifies the information regarding a parameter contained in measurement outcomes, indicating how well a quantum state can be distinguished from nearby states concerning the measured parameter [40].
Another utilization of QFI involves employing it as a measure to quantify quantum correlations. Hence, Girolami et al. [40] showed that local quantum Fisher information (LQFI), a computable measure of discord-type quantum correlations, can be expressed as follows
@@ V LQFI of gravcats
This discord-type measure (denoted by $\mathcal{F}$) builds upon QFI ($F$) and represents the optimal LQFI when $H_{A}$ (a measurement operator) is applied to subsystem $A$ belonging to bipartite system $AB$.
@@ VII Comparative analysis of quantum resources
For $\mu=1$, as depicted in Fig. 8, all quantities remain constant over time because the density operator (9) is time-independent in this special case. This finding underscores the critical role of classical correlations in preserving quantum resources. Conversely, for $\mu<1$, exponential decay is observed in all quantities. Lower values of $\mu$ correspond to more pronounced decay, highlighting the detrimental effect of reduced classical correlations on the preservation of quantum resources. In the extreme case of $\mu=0$, quantum coherence and LQFI asymptotically approach zero, while the Bell parameter converges to $2$, indicating the loss of non-classical correlations.
@@
For a more detailed examination of the relationship between the classical correlation parameter ($\mu$) and various quantum resources – quantum coherence, LQFI, and Bell non-locality parameter, we draw Fig. 10. This figure consists of three plots, each representing a different quantum resource. In each plot, there are five curves, corresponding to different time instants [by fixing $f$ according to Eqs. (7) and (8)], with the bottom curve representing the infinite-time limit (i.e., $f=0$).
The overall trend across all plots indicates that quantum resources generally increase with increasing values of $\mu$. This confirms that higher classical correlations contribute to the preservation of quantum resources. Importantly, in each plot, curves corresponding to earlier times lie above those for later times, demonstrating that as time progresses, quantum resources generally decrease due to decoherence. However, a critical observation is that as $\mu$ approaches $1$, these curves converge, indicating that strong classical correlations can mitigate the effects of decoherence, even for the infinite-time limit (see also Refs. [30, 48, 49]).
Focusing specifically on Fig. 10 (c), we see that as $\mu$ approaches $1$, the Bell parameter converges to a limit close to $2\sqrt{2}$, which is the maximum possible value of non-locality. This is significant because it suggests that even after infinite time, with sufficient classical correlations, quantum non-locality can be preserved, indicating a resistance to complete degradation.
The physical interpretation of the results provided in Fig. 10 indicates that when classical correlations are maximized ($\mu\rightarrow 1$), quantum resources are preserved. This robustness arises because classical correlations effectively synchronize the environment’s effect on the system, reducing the differential impact of decoherence on quantum resources [30]. The fact that these findings hold true even in the infinite-time limit highlights the importance of classical correlations in quantum information processing and suggests that maintaining high classical correlations can be a strategy to ensure long-term quantum resource preservation.
As we conclude this section, it is important to address further observations regarding the discussed phenomena. Firstly, the classical correlations between successive channel actions serve to enhance the overall coherence of both qubits collectively, without affecting the coherence of the single qubit. Secondly, we identify two potential sources of memory effects in the dephasing channel under scrutiny. The first originates from temporal correlations inherent in the time evolution of each qubit, while the second arises from classical correlations between successive channel actions [50]. These factors collectively influence the decoherence process of the system. Notably, the damped oscillations of quantum resources are attributed to the first type of non-Markovian memory effects, while their increase at a fixed time is attributed to the second type of memory effects [30].